OpenStudy (anonymous):

Stuck on some simple algebra in the very first example of the first lecture (talk about rusty!). I understand the algebra all the way until the end where: (1/Δx)[(x0-(x0+Δx))/((x0+Δx)x0)] = (1/Δx)[(-Δx/((x0+Δx)x0)] How does the numerator reduce to -Δx? Wouldn't x0-(x0+Δx) = Δx?

OpenStudy (anonymous):

After some thought, I believe I answered my own question. The minus sign is transitive and is reduced to x0 - x0 - Δx.

OpenStudy (phi):

you "distribute" the minus sign -1*(a +b) = -1*a + -1*b = -a - b